# Interlayer interactions in $\text{La}_{3}\text{Ni}_{2}\text{O}_{7}$ under pressure: from $s^{\pm}$ to $d_{xy}$-wave superconductivity

### v1

@@ Abstract
As discussed in Sec. Superconducting symmetries of the main text, the role of considering a two-orbital model for the superconductivity of bilayer nickelates can be tested by solving the eigenvalue Eq. (7) for the desired pockets only. In the argumentation, we note that we have checked that no individual pocket shows the desired $s^{\pm}$- and $d_{xy}$-wave interplay as the two-orbital model. The calculations are shown in Fig. S1. Panels (a) and (b) show no change in the $d_{xy}$ symmetry of the $\beta$ pocket as the interlayer coupling varies from $U_{\perp}/U=0$ (a) to $U_{\perp}/U=0.3$ (b). In fact, the interlayer interactions do not change the symmetry in any three cases explored. As shown in panel (c), a $d_{x^{2}-y^{2}}$-wave dominates pairing of individual $\gamma$ pocket, and a $g$-wave symmetry is obtained for the $\alpha$ pocket alone (d).
Figure S1: Same as Fig. 3 of the main text but for individual pockets $\beta$ (a,b), $\gamma$ (c), and $\alpha$ (d).
Figure S2: Cuts of the phase diagram shown in Fig. 2 in the main text.

### v3

@@ Appendix B Pairing symmetry of individual pockets
Fig. 6 shows results for couples of pockets but we have also checked that conclusions hold for the individual pockets as well. In Fig. 9, panels (a) and (b) show no change in the $d_{xy}$ symmetry of the $\beta$ pocket as the interlayer coupling varies from $U_{\perp}/U=0$ (a) to $U_{\perp}/U=0.3$ (b). In fact, the interlayer interactions do not change the symmetry in any three cases explored. As shown in panel (c), a $d_{x^{2}-y^{2}}$-wave dominates pairing of individual $\gamma$ pocket, and a $g$-wave symmetry is obtained for the $\alpha$ pocket alone (d). None of the single-pocket solutions reproduces the $s^{\pm}\rightarrow d_{xy}$ transition as a function of interlayer interactions.